Week 4, Question 2: Use the force method to determine the force in each member of the truss. Take a = 7 kN, b = 8 kN, c = 11e) FcD = f) Fce = g) FCF = h) FDE = i) FDF = j) FEF =part a, b, c & d are correct
Answer #1

• Statical element or indetene maaj +3 - 10-2x6+3 - 13-12 IV No.of in Detmery = 1 Remove member DE Ra AL RARE = 7+3+1 = 26 KH(Facto] TKN FOD tomos 3 dlm inė 371w 57 Efy=0 7+11-FBc sing =D foca e fbc= 4513 = 2513 KN teresting FBD = Foc coso = 2513x3 1ref= 2513 KN tensim Efn=0 FBc Coso - FCE + FefCOSO O IFCE=0 Joint E 4 FEF Fef=\SKN compressin 15 DE Now and applied Remore un[RE =o) [Ran=0 RA=o S - Joint a KAB=0 Kaczo - - - Joint I D Kof Icoso = KDF > Koop compressin Joint a keo= kefsino tef= 2 2 2Joint KEF= 1 KCE = 1/2 KCE Scanned with CamScannerBC 7. Member Force (P Unit Force(Kg AB T-110 ACO 0 BD -6 O 7.2111 CD -8 -0.5547002 CE 0 -0.83205029 | 7.2111 DF -0.83205029 E2- - PleL AE - 66.4931 12.5957 3 k2L AE 1x=-51279025 Force - x in all member TE=p+kx FAB=-11 KN I FBC= I T-BC= 7.2111 KN 7.21​​​​​​